# Entropy Stability and Spectral Concentration under Convex Block Constraints

### v1

@@ 4.5 Conditional entropy stability
Lemma 4.4 (Conditional entropy stability).
Let $\mathcal{C}_{i}\subset\mathcal{S}(\mathcal{H}_{i})$ be compact and convex, and let $\sigma_{i}$ minimize entropy on $\mathcal{C}_{i}$. Then for all $\rho_{i}\in\mathcal{C}_{i}$,
$S(\rho_{i})-S(\sigma_{i})\;\geq\;\tfrac{1}{2}\|\rho_{i}-\sigma_{i}\|_{1}^{2}.$
Proof.
Since $\sigma_{i}$ is an entropy minimizer on $\mathcal{C}_{i}$, we may assume $\operatorname{supp}(\rho_{i})\subseteq\operatorname{supp}(\sigma_{i})$; otherwise the relative entropy is infinite and the inequality holds trivially. In this case,
$S(\rho_{i})-S(\sigma_{i})=D(\rho_{i}\|\sigma_{i}),$
and Pinsker’s inequality yields
$D(\rho_{i}\|\sigma_{i})\geq\tfrac{1}{2}\|\rho_{i}-\sigma_{i}\|_{1}^{2}.$
@@ 4.6 Assembly of the entropy estimate
Applying Lemmas 4.3 and 4.4 to the marginal and conditional contributions, respectively, we obtain the quantitative lower bound
@@ 4.7 Blockwise trace norm control
Squaring and applying $(a+b)^{2}\leq 2a^{2}+2b^{2}$ yields the claim. ∎
@@ 4.8 Conclusion of the proof
This completes the proof of Theorem 4.1. ∎
@@ 5 Sharpness and Optimality of the Stability Exponent
All examples are finite-dimensional and require no additional structure beyond that developed in Sections 3–4.
@@ 5.1 The classical simplex model
Since $H$ is strictly concave on the face $F$, the quadratic form is negative definite, implying the stated asymptotic behavior. ∎
@@ 5.2 Quantum block-diagonal models
Minimization forces $p=q$ extreme and each $S(\rho_{i})=0$, hence $\rho_{i}$ pure. ∎
@@ 5.3 Sharpness under block perturbations
since blocks are orthogonal. The claim follows directly from Proposition 5.1. ∎
@@ 5.4 Failure of linear stability
Under the hypotheses of Theorem 4.1, the exponent $2$ is optimal. In general, there exists no constant $C>0$ such that
$S(\rho)-S_{\min}\geq C\,\operatorname{dist}_{1}(\rho,\mathcal{M})$
@@
which precludes any linear lower bound. ∎
@@ 5.5 Interpretation
The obstruction to linear stability arises from flat directions along faces of the marginal polytope $\Pi$. Entropy exhibits genuine curvature only transverse to these faces, leading to quadratic — but not linear — rigidity.
@@
6 A Concrete Application: Stability of Constrained Gibbs Ensembles
In this section we present a fully explicit application of Theorem 4.1. The purpose is twofold: first, to demonstrate that the stability theorem yields genuinely new quantitative information in a natural physical setting; second, to clarify the role and optimality of the $\sqrt{\varepsilon}$ stability scale.
6.1 Block constraints from conserved quantities
Let $\mathcal{H}$ be a finite-dimensional Hilbert space and let $H_{0}$ be a self-adjoint operator with discrete spectrum
$H_{0}=\sum_{i=1}^{r}E_{i}P_{i},$
where $\{P_{i}\}$ are mutually orthogonal spectral projections. We assume that the eigenspaces $\mathcal{H}_{i}:=P_{i}\mathcal{H}$ are nontrivial.
Such a decomposition arises whenever a system possesses a conserved quantity, or when a Hamiltonian is measured with finite resolution.
@@ 6.1 Block constraints from conserved quantities
6.2 Entropy minimization under energy constraints
Fix a probability vector $q=(q_{1},\dots,q_{r})$ with $q_{i}>0$ and $\sum_{i}q_{i}=1$. Consider the affine constraint set
$\mathcal{C}_{q}:=\left\{\rho\in\mathcal{S}_{\mathrm{bd}}(\mathcal{H}):\operatorname{Tr}(P_{i}\rho)=q_{i}\;\;\forall i\right\}.$
This set consists of all block-diagonal states with fixed energy-sector populations. It is block-convex in the sense of Definition 3.3, with
$\Pi=\{q\},\qquad\mathcal{C}_{i}=\mathcal{S}(\mathcal{H}_{i}).$
@@ 6.2 Entropy minimization under energy constraints
where each $|\psi_{i}\rangle$ is an arbitrary unit vector in $\mathcal{H}_{i}$. Equivalently, the minimizer is unique up to unitary transformations within each energy block.
Proof.
For $\rho=\oplus_{i}q_{i}\rho_{i}\in\mathcal{C}_{q}$, entropy decomposes as
@@
Since $H(q)$ is constant on $\mathcal{C}_{q}$, entropy is minimized by minimizing each $S(\rho_{i})$, which forces $\rho_{i}$ to be pure. ∎
6.3 Quantitative stability of near-minimal entropy states
We now apply Theorem 4.1 in this setting.
Theorem 6.2 (Stability of constrained Gibbs ensembles).
There exists a constant $C>0$, depending only on the probability vector $q$ (and in particular independent of the dimensions $\dim\mathcal{H}_{i}$), such that for all $\rho\in\mathcal{C}_{q}$,
$S(\rho)-S_{\min}\;\geq\;C\,\|\rho-\sigma_{q}\|_{1}^{2},$
where $S_{\min}=S(\sigma_{q})=H(q)$.
Proof.
The constraint set $\mathcal{C}_{q}$ is block-convex with a single marginal point $q$. The claim follows directly from Theorem 4.1. ∎
@@ 6.4 Explicit constant in the uniform case
The stability constant can be computed explicitly in a canonical case.
Assume that $q_{i}=\frac{1}{r}$ for all $i$, i.e. the uniform energy distribution. Then
@@
Proposition 6.3 (Explicit stability constant).
If $q$ is uniform, then
$S(\rho)-S_{\min}\;\geq\;\frac{1}{2r}\,\|\rho-\sigma_{q}\|_{1}^{2}\qquad\forall\rho\in\mathcal{C}_{q}.$
Proof.
Since the marginal distribution is fixed, entropy excess arises solely from internal block entropy:
$S(\rho)-S_{\min}=\frac{1}{r}\sum_{i=1}^{r}S(\rho_{i}).$
By Pinsker’s inequality applied in each block,
$S(\rho_{i})\geq\tfrac{1}{2}\|\rho_{i}-|\psi_{i}\rangle\langle\psi_{i}|\|_{1}^{2}.$
Summing over $i$ and using blockwise trace-norm additivity yields the claim. ∎
6.5 Stability at scale $\varepsilon$
The previous inequality implies
$\|\rho-\sigma_{q}\|_{1}\;\leq\;\sqrt{2r\,(S(\rho)-S_{\min})}.$
Thus an entropy defect of order $\varepsilon$ forces the state to be within $O(\sqrt{\varepsilon})$ trace distance of the extremal Gibbs ensemble. By Section 5, the quadratic exponent is optimal.
In particular, no linear bound of the form
@@ 6.5 Stability at scale $\varepsilon$
can hold uniformly, even in this elementary physical model.
6.6 Interpretation
The result provides a quantitative version of the heuristic principle that near-minimal entropy under fixed macroscopic constraints forces microscopic proximity to an extremal equilibrium configuration. Unlike classical treatments, this conclusion is obtained without invoking dynamics, spectral gaps, or Markovian assumptions.
6.7 Why this application is nontrivial
This result is not a consequence of standard entropy inequalities:
  • Pinsker’s inequality requires a fixed reference state, whereas here the reference state emerges from constrained minimization.
  • Log-Sobolev inequalities require a Markov semigroup and spectral gap, which are absent in the present static setting.
  • The von Neumann entropy is not globally strongly convex; strong convexity emerges only after restriction to the block-convex constraint.
Accordingly, the stability bound obtained here provides genuinely new quantitative information, even in finite-dimensional Gibbs-type models.
@@ 7 Relation to Classical Entropy Inequalities
The stability inequality established in Theorem 4.1 interacts with, but is fundamentally distinct from, several well-known entropy inequalities in information theory and mathematical physics. We clarify these relationships in this section.
7.1 Pinsker and Csiszár–Kullback inequalities
Let $\rho,\sigma\in\mathcal{S}(\mathcal{H})$ with $\operatorname{supp}(\rho)\subseteq\operatorname{supp}(\sigma)$. The classical Pinsker inequality states
$S(\rho\|\sigma)\;\geq\;\tfrac{1}{2}\|\rho-\sigma\|_{1}^{2},$
where $S(\rho\|\sigma)=\operatorname{Tr}(\rho(\log\rho-\log\sigma))$ is the relative entropy.
This inequality is fundamentally two-state in nature: it compares a state $\rho$ to a fixed reference state $\sigma$.

### v2

@@ 4.5 Conditional entropy stability
Definition 4.6 (Nondegenerate constrained minimizer).
Let $C\subset S(H)$ be a constraint set and $\sigma\in C$. We say that $\sigma$ is a nondegenerate constrained minimizer of f:=Sf:=-S on the fixed-support stratum if there exists $m>0$ such that, for all admissible tangent directions $X$ to $C\cap D(PH)$ at $\sigma$ (with $P=\operatorname{supp}(\sigma)$), one has
$\langle X,\nabla^{2}f(\sigma)[X]\rangle\geq m\|X\|_{2}^{2}.$
where $\|\cdot\|_{2}$ denotes the Hilbert–Schmidt norm.
Lemma 4.7 (Entropy stability under confining constraints).
Let $\mathcal{C}$ be a convex constraint set of quantum states on a finite-dimensional Hilbert space $\mathcal{H}$, and let $\sigma\in\mathcal{C}$ be an entropy minimizer.
Assume that there exists a neighbourhood $U$ of $\sigma$ such that all admissible perturbations preserve the support of $\sigma$, i.e.
$P(\rho-\sigma)Q=Q(\rho-\sigma)P=0\qquad\text{for all }\rho\in\mathcal{C}\cap U,$
where $P$ is the support projection of $\sigma$ and Q=IPQ=I-P.
Assume moreover that $\sigma$ is a nondegenerate constrained minimizer of S-S on this fixed-support stratum.
Then there exist constants $c>0$ and $\varepsilon_{0}>0$ such that for all $\rho\in\mathcal{C}$ with
$S(\rho)-S(\sigma)\leq\varepsilon_{0},$
one has
$S(\rho)-S(\sigma)\ \geq\ c\,\|\rho-\sigma\|_{1}^{2}.$
Equivalently,
$\|\rho-\sigma\|_{1}\ \leq\ C\,\sqrt{S(\rho)-S(\sigma)},$
for some constant $C>0$.
Remark 4.8 (Scope of Lemma 4.7).
Lemma 4.7 applies to the fixed-support (confining) regime, that is, to situations in which admissible perturbations of the constrained entropy minimizer remain within a fixed-rank support stratum.
Proof 4.9.
By the confining (fixed-support) hypothesis, there exists a neighbourhood $U$ of $\sigma$ such that every admissible $\rho\in\mathcal{C}\cap U$ satisfies
$Q\rho=\rho Q=0,\qquad P=\operatorname{supp}(\sigma),\quad Q=I-P.$
Hence all admissible states lie in the fixed-support stratum
$\mathcal{S}(P\mathcal{H}):=\{\rho\in\mathcal{S}(\mathcal{H}):Q\rho=\rho Q=0\},$
on which $\sigma$ is faithful, i.e. $P\sigma P\succ 0$ as an operator on $P\mathcal{H}$.
On this stratum the von Neumann entropy is $C^{2}$ in a neighbourhood of $\sigma_{i}$, and f:=Sf:=-S has a nondegenerate constrained minimum at $\sigma_{i}$ by assumption. Therefore, in a local chart of the admissible set $C_{i}\cap U$ around $\sigma_{i}$, the restricted Hessian satisfies
$\nabla^{2}f(\sigma_{i})\ \succeq\ m_{i}\,I$
on the tangent space of $C_{i}$ at $\sigma_{i}$, for some $m_{i}>0$. A second-order Taylor expansion then yields, for $\rho_{i}$ sufficiently close to $\sigma_{i}$,
$f(\rho_{i})-f(\sigma_{i})\ \geq\ \frac{m_{i}}{2}\,\|\rho_{i}-\sigma_{i}\|_{2}^{2},$
and hence
$S(\rho_{i})-S(\sigma_{i})=f(\sigma_{i})-f(\rho_{i})\ \geq\ \frac{m_{i}}{2}\,\|\rho_{i}-\sigma_{i}\|_{2}^{2}.$
Finally, using $\|X\|_{2}\geq\|X\|_{1}/\sqrt{\dim(P_{i}\mathcal{H}_{i})}$, we obtain
$S(\rho_{i})-S(\sigma_{i})\ \geq\ \frac{m_{i}}{2\,\dim(P_{i}\mathcal{H}_{i})}\,\|\rho_{i}-\sigma_{i}\|_{1}^{2},$
which proves the claimed quadratic stability bound. Since the support projection $P$ and the admissible block dimensions are fixed by the constraint geometry, the resulting constants depend only on the constraint set and not on the ambient Hilbert space dimension.
Remark 4.10 (Why this excludes the boundary–transversal regime).
The argument above crucially uses the fixed-support (confining) hypothesis. In the boundary–transversal regime, admissible perturbations necessarily generate support–kernel coupling, and the entropy variation acquires an unavoidable $\delta^{2}\log(1/\delta)$ correction.
@@ 4.6 Assembly of the entropy estimate
Applying Lemmas 4.3 and 4.7 to the marginal and conditional contributions, respectively, we obtain the quantitative lower bound
@@ 4.7 Blockwise trace norm control
Squaring and applying $(a+b)^{2}\leq 2a^{2}+2b^{2}$ yields the claim.
@@ 4.8 Conclusion of the proof
This section makes explicit the dependence of the stability constant $C=C(\Pi,\{\mathcal{C}_{i}\})$ appearing in Theorem 4.1, by expressing it in terms of the marginal and conditional curvature constants introduced in the preceding lemmas.
@@
Since $M$ is compact and the confining/nondegeneracy hypotheses hold uniformly along $M$ by assumption, the local constants produced by Lemmas 4.34.7 admit a strictly positive global minimum over a finite subcover of $M$.
This completes the proof of Theorem 4.1.
@@ 5 Sharpness and Optimality of the Stability Exponent
All examples are finite-dimensional and require no additional structure beyond that developed in Sections 3-4.
@@ 5.1 The classical simplex model
Since $H$ is strictly concave on the relative interior of the face $F$, the Hessian of $H$ is negative definite on the tangent space of $F$ at $q$, implying the stated asymptotic behavior.
@@ 5.2 Quantum block-diagonal models
Minimization forces $p=q$ extreme and each $S(\rho_{i})=0$, hence $\rho_{i}$ pure.
@@ 5.3 Sharpness under block perturbations
since blocks are orthogonal. The claim follows directly from Proposition 5.1.
@@ 5.4 Failure of linear stability
Under the hypotheses of Theorem 4.1, the exponent $2$ is optimal. In general, there exists no constant $C>0$ and no exponent $\alpha<2$ such that
$S(\rho)-S_{\min}\geq C\,\operatorname{dist}_{1}(\rho,\mathcal{M})^{\alpha}$
@@
which precludes any linear lower bound.
@@ 5.5 Interpretation
The obstruction to linear stability arises from flat directions along faces of the marginal polytope $\Pi$. Entropy exhibits genuine curvature only transverse to these faces, leading to quadratic - but not linear - rigidity.
@@
6 Example: Pure State Entropy Minimization
In this section we provide a concrete example where the stability constants in Theorem 4.1 can be computed explicitly. The constraint set fixes block populations while allowing states within each block to vary freely, yielding entropy minimizers that are pure states in each block. This example illustrates the decomposition of the stability constant into marginal and conditional components and demonstrates the sharpness of the quadratic bound.
6.1 Setup
Let $\mathcal{H}$ be a finite-dimensional Hilbert space and suppose
$\mathcal{H}=\bigoplus_{i=1}^{r}\mathcal{H}_{i}$
is a fixed orthogonal decomposition into nontrivial subspaces, with $P_{i}$ denoting the orthogonal projection onto $\mathcal{H}_{i}$.
@@ 6.1 Setup
Such decompositions arise naturally from symmetry reductions, superselection rules, or coarse-grained descriptions of larger systems.
6.2 Fixed Block Populations
Fix a probability vector $q=(q_{1},\ldots,q_{r})$ with $q_{i}>0$ and $\sum_{i}q_{i}=1$. Consider the affine constraint set
$\mathcal{C}_{q}:=\left\{\rho\in\mathcal{S}_{\mathrm{bd}}(\mathcal{H}):\mathrm{Tr}(P_{i}\rho)=q_{i}\text{ for all }i\right\}.$
This constraint fixes the population (total probability) in each block while allowing the internal state within each block to vary. The set $\mathcal{C}_{q}$ is block-convex in the sense of Definition 3.3, with $\Pi=\{q\}$ (a single point) and $\mathcal{C}_{i}=\mathcal{S}(\mathcal{H}_{i})$ (the full state space of block $i$).
@@ 6.2 Fixed Block Populations
where each $|\psi_{i}\rangle$ is an arbitrary unit vector in $\mathcal{H}_{i}$. Equivalently, the minimizers are product states with pure states in each block, unique up to unitary transformations within each block.
Proof 6.2.
For $\rho=\bigoplus_{i}q_{i}\rho_{i}\in\mathcal{C}_{q}$, entropy decomposes as
@@
Since $H(q)$ is constant on $\mathcal{C}_{q}$, entropy is minimized by minimizing each $S(\rho_{i})$. The von Neumann entropy $S(\rho_{i})$ is minimized when $\rho_{i}$ is a pure state (rank-one projection), giving $S(\rho_{i})=0$.
6.3 Quantitative Stability Estimate
We now derive an explicit quantitative stability estimate in this setting.
Theorem 6.3 (Stability for fixed block populations).
Let $q=(q_{1},\dots,q_{r})$ with $q_{i}>0$ and $\sum_{i}q_{i}=1$, and let $\mathcal{C}_{q}$ be as above. Let
$\mathcal{M}_{q}:=\left\{\bigoplus_{i=1}^{r}q_{i}\,|\psi_{i}\rangle\langle\psi_{i}|:\ |\psi_{i}\rangle\in\mathcal{H}_{i},\ \|\psi_{i}\|=1\right\}$
denote the set of entropy minimizers in $\mathcal{C}_{q}$, and write $\mathrm{dist}_{1}(\rho,\mathcal{M}_{q}):=\inf_{\sigma\in\mathcal{M}_{q}}\|\rho-\sigma\|_{1}$.
Then for every $\rho=\bigoplus_{i}q_{i}\rho_{i}\in\mathcal{C}_{q}$, the entropy gap satisfies
$S(\rho)-S_{\min}=\sum_{i=1}^{r}q_{i}\,S(\rho_{i}),\qquad S_{\min}=H(q),$
and the following quantitative implication holds: if $\mathrm{dist}_{1}(\rho,\mathcal{M}_{q})\leq 1$, then
$S(\rho)-S_{\min}\;\geq\;\frac{1}{2}\,\mathrm{dist}_{1}(\rho,\mathcal{M}_{q})^{2}.$
Equivalently, whenever $S(\rho)-S_{\min}\leq\frac{1}{2}$, one has
$\mathrm{dist}_{1}(\rho,\mathcal{M}_{q})\;\leq\;\sqrt{2\,(S(\rho)-S_{\min})}.$
Proof 6.4.
The entropy decomposition follows from block-diagonality: for $\rho=\bigoplus_{i}q_{i}\rho_{i}$, one has $S(\rho)=H(q)+\sum_{i}q_{i}S(\rho_{i})$.
For each block $i$, define the distance to the set of pure states:
$d_{i}:=\mathrm{dist}_{1}\bigl(\rho_{i},\mathrm{Pure}(\mathcal{H}_{i})\bigr)=\inf_{\|\psi\|=1}\|\rho_{i}-|\psi\rangle\langle\psi|\|_{1}.$
By trace-norm additivity over orthogonal blocks,
(2) $\mathrm{dist}_{1}(\rho,\mathcal{M}_{q})=\sum_{i=1}^{r}q_{i}\,d_{i}.$
Let $\lambda_{i}:=\lambda_{\max}(\rho_{i})$ be the largest eigenvalue of $\rho_{i}$. Choosing $|\psi_{i}\rangle$ to be an eigenvector corresponding to $\lambda_{i}$ gives $d_{i}=2(1-\lambda_{i})$.
By Schur concavity of entropy, among density matrices with fixed largest eigenvalue $\lambda_{i}$, the entropy $S(\rho_{i})$ is minimized when the remaining probability mass $1-\lambda_{i}$ is concentrated on a single orthogonal eigenvector. This gives the lower bound
$S(\rho_{i})\ \geq\ h(1-\lambda_{i}),\qquad h(t):=-(1-t)\log(1-t)-t\log t,$
where $h$ is the binary entropy function.
If $d_{i}\leq 1$, then $t_{i}:=1-\lambda_{i}=d_{i}/2\in[0,\frac{1}{2}]$. On this interval, $h$ satisfies the quadratic lower bound
[…]
which follows from $h(0)=h^{\prime}(0)=0$ and $h^{\prime\prime}(t)=\frac{1}{t(1-t)}\geq 4$ on $0,\frac{1}{2}]$.
Therefore, whenever $d_{i}\leq 1$,
$S(\rho_{i})\ \geq\ \frac{1}{2}\,d_{i}^{2}.$
Assuming $\mathrm{dist}_{1}(\rho,\mathcal{M}_{q})\leq 1$, we have $\sum_{i}q_{i}d_{i}\leq 1$, which implies $d_{i}\leq 1$ for all $i$. Summing the previous inequality with weights $q_{i}$ yields
$S(\rho)-S_{\min}=\sum_{i}q_{i}S(\rho_{i})\ \geq\ \frac{1}{2}\sum_{i}q_{i}d_{i}^{2}\ \geq\ \frac{1}{2}\left(\sum_{i}q_{i}d_{i}\right)^{2},$
where the last step uses Jensen’s inequality for the convex function $x\mapsto x^{2}$. Using (2) gives the claimed bound. The final implication is immediate.
@@ 6.4 Explicit Constant in the Uniform Case
The stability constant can be computed explicitly when all block populations are equal.
Assume that $q_{i}=\frac{1}{r}$ for all $i$. Then the entropy minimizers are
@@
Corollary 6.5 (Uniform block populations).
Assume that $q_{i}=\frac{1}{r}$ for all $i$. Then for all $\rho\in\mathcal{C}_{q}$ with $\mathrm{dist}_{1}(\rho,\mathcal{M}_{q})\leq 1$,
$S(\rho)-S_{\min}\;\geq\;\frac{1}{2}\,\mathrm{dist}_{1}(\rho,\mathcal{M}_{q})^{2}.$
In particular, if $S(\rho)-S_{\min}\leq\frac{1}{2}$, then
$\mathrm{dist}_{1}(\rho,\mathcal{M}_{q})\;\leq\;\sqrt{2\,(S(\rho)-S_{\min})}.$
Proof 6.6.
This follows immediately from Theorem 6.3 specialized to the uniform marginal distribution $q$.
6.5 Optimality of the Quadratic Exponent
The inequality in Theorem 6.3 gives
$\mathrm{dist}_{1}(\rho,\mathcal{M}_{q})\;\leq\;\sqrt{2\,(S(\rho)-S_{\min})}.$
Thus, an entropy gap of order $\varepsilon$ implies the state is within $O(\sqrt{\varepsilon})$ in trace distance from the set of entropy minimizers. By Section 5, this quadratic exponent is optimal: no linear bound of the form
@@ 6.5 Optimality of the Quadratic Exponent
can hold uniformly, even in this simple example.
This demonstrates the sharpness of the quadratic stability principle and confirms that the $\sqrt{\varepsilon}$ distance scaling cannot be improved without additional structure.
6.6 Relation to Standard Entropy Inequalities
This stability result is not a direct consequence of standard entropy inequalities:
  • Pinsker’s inequality $D(\rho\|\sigma)\geq\frac{1}{2}\|\rho-\sigma\|_{1}^{2}$ provides a lower bound relating relative entropy to trace distance. In our setting, the reference state $\sigma_{q}$ emerges from constrained optimization rather than being fixed a priori. The constraint structure (fixed block populations) enables explicit computation of the stability constant.
  • Entropy continuity bounds (Fannes, Audenaert, Winter) provide global moduli of continuity for von Neumann entropy. Our result is a local stability estimate near constrained minimizers, exploiting the specific geometry of block-separable constraints to obtain sharp constants.
  • Strong convexity: The von Neumann entropy is globally concave but not strongly concave. Quadratic stability emerges after restriction to the constraint set $\mathcal{C}_{q}$ and localization near minimizers in the fixed-support regime.
The explicit computation of the stability constant $C=\frac{1}{2}$ in Theorem 6.3 provides quantitative information beyond what standard entropy inequalities yield for this constrained optimization problem.
@@ 7 Relation to Classical Entropy Inequalities
The stability inequality established in Theorem 4.1 interacts with, but is fundamentally distinct from, several well-known entropy inequalities in information theory and convex optimization. We briefly clarify these relationships and emphasize the nonstandard features of the present result.
7.1 Relation to Pinsker, continuity, and functional inequalities
Pinsker’s inequality [Pinsker1964, Csiszar1967] provides a quantitative relation between relative entropy and trace distance by comparing two fixed states. Related continuity bounds for the von Neumann entropy were established by Fannes [Fannes1973], Audenaert [Audenaert2007], and Winter [Winter2016], yielding sharp global moduli of continuity for entropy as a function of states.